# APB: Accelerating Distributed Long-Context Inference by Passing Compressed Context Blocks across GPUs

### v1

@@ Abstract
Yuxiang Huang1††thanks: indicates equal contribution., Mingye Li2∗††thanks: Work done during internship at TsinghuaNLP., Xu Han1††thanks: indicates corresponding authors., Chaojun Xiao1, Weilin Zhao1, Ao Sun3, Hao Zhou4, Jie Zhou4, Zhiyuan Liu1, Maosong Sun1 1NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China. 2Department of CS&T, Central South University, Changsha, China. 3BUPT, Beijing, China. 4Pattern Recognition Center, WeChat Al, Tencent Inc. huang-yx21@mails.tsinghua.edu.cn, lmy2004@csu.edu.cn, han-xu@tsinghua.edu.cn, xiaocj20@mails.tsinghua.edu.cn
@@ B.1 Training of the Compressor
We use the retraining heads proposed in Huang et al. (2024) as our compressor. Following Huang et al. (2024), we set the intermediate size of the retaining heads to 1024. We utilize the first 3000 samples from the LongAlign (Bai et al., 2024) dataset to generate the training labels, and we train the retaining heads with the a frozen model backbone for 3000 steps in 1 epoch. The batch size set is 1 and the maximum input length is set to 10240. We use AdamW as the optimizer and the learning rate is set to 5e-4, with set to 0.9 and set to 0.95. We apply a linear scheduler whose number of warmup steps is set to 300. Notably, the training loss of the retaining head consists of a regression loss and a smoothing loss, following the setting of Huang et al. (2024), and the balance factor is set to 0.0025. The gradient clipping value is set to 0.5 to avoid gradient explosion.
@@ Appendix D Task Abbreviations
Here, we provide a mapping between the benchmark task abbreviations and their full names used in the main text in Table 7 and 8.
@@ E.2 Hyperparameter Stability
E.3 The Detailed Experimantal Results
Here, we present the detailed experimental results to extensively illustrate each figure in the main text, Section E.1 and E.2. In Section 4.2, the detailed experimental results of Figure 3 are provided in Table 9 and Table 12. Table 9 provides the accurate speed for all methods on Bench, while the speed is the average value across all tasks. By contrast, Table 12 reports the precise speed measurements for all methods on RULER, with the values similarly averaged over all tasks. The performance of each method in Figure 3 is provided in Table 1 and Table 2. In Section 4.3, Figure 14 and Figure 15 present the detailed results corresponding to Figure 4. Specifically, Table 14 aligns with Figure 44 (a) while Table 15 corresponds to Figure 44 (b). In Section 4.4, the detailed experimental results of Figure 5 are provided in Table 13. In Section E.1, we add the breakdown analysis of the prefill and decoding wall-time for various methods, while the results are illustrated in Figure 6. The detailed experimental results are provided in Table 10. In Section 1, the detailed experimental results of Figure 1 are provided in Table 11.

### v2

@@ Abstract
Yuxiang Huang1††thanks: indicates equal contribution., Mingye Li2∗††thanks: Work done during internship at TsinghuaNLP., Xu Han1††thanks: indicates corresponding authors., Chaojun Xiao1, Weilin Zhao1, Ao Sun3, Hao Zhou4, Jie Zhou4, Zhiyuan Liu1, Maosong Sun1 1NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China. 2Department of CS&T, Central South University, Changsha, China. 3BUPT, Beijing, China. 4Pattern Recognition Center, WeChat AI, Tencent Inc. huang-yx21@mails.tsinghua.edu.cn, lmy2004@csu.edu.cn, han-xu@tsinghua.edu.cn, xiaocj20@mails.tsinghua.edu.cn
@@ 4.4 Ablation Studies
Short-Context Performance. While APB exhibits outstanding performance and inference speed on long-context benchmarks, its effectiveness and efficiency on short-context inputs (i.e., 4K tokens) are equally important. To evaluate this, we conduct experiments on RULER using 4K input lengths, reporting both task performance and inference speed. We compare APB against FlashAttn using the Llama-3-8B-instruct-1M model, and present the results in Table 5.
The results demonstrate that APB consistently achieves superior speed and performance on short-context inputs. Even in such scenarios, APB outperforms FullAttn without any degradation in performance or slowdown in inference. Please refer to Appendix E.3 for more details.
| Method | RULER-4K | Speed(tok/s). |
| FlashAttn | 94.54 | 6247.58 |
| APB | 94.89 | 6597.47 |
Orthogonality to Quantization. APB exhibits strong orthogonality and can be integrated with other compression methods, such as KV quantization. To demonstrate this, we tested APB in combination with KV cache quantization techniques. We integrate the KV cache quantization method officially supported by Hugging Face Transformers222https://github.com/huggingface/transformers into the APB framework. This method is inspired by KIVI (Liu et al., 2024) and implemented using the HQQ backend333https://github.com/mobiusml/hqq. We evaluate the integration on the RULER benchmark using 50 entries per task with LLaMA-3.1-8B-Instruct as the backbone. The KV cache is quantized to 4 bits. For comparison, the results of FullAttn, MInference, StarAttn, and standalone APB are taken from Table 2, which used 500 entries per task. The results are summarized in Table 6.
The results shows that APB is compatible with KV quantization. In this setting, the stadalone APB shows a minor performance drop compared with FullAttn, and APB+HQQ-4bits also shows a small performance degredation compared with standalone HQQ-4bits quantization. However, it is worth noticing that APB+HQQ-4bits can still ourperform the baselines (MInference and StarAttn), as the baselines do not use quantization. This setting is able to reduce the size of KV cache to 4 and enhance prefill speed at the meantime, showing a promising balance of task performance and inference efficiency.
| Method | Original (M) | Quantized (M-4bits) |
| FullAttn | 82.20 | 81.73 |
| MInference | 72.97 | - |
| StarAttn | 76.84 | - |
| APB | 81.63 | 79.13 |
Training Generalizability. To demonstrate that the training of the retaining heads is robust and generalizable, we conduct an ablation study on their training data. Specifically, we train the retaining heads on datasets including LongAlpaca (Chen et al., 2024), LongAlign, and AntiHaystack444https://huggingface.co/datasets/wenbopan/anti-haystack, using Llama-3.1-8B-instruct as the backbone model. We then evaluate APB with these trained retaining heads on RULER across 20 samples per task. Other settings are aligned to Table 2. The results are presented in Table 7.
The similar results across different training data indicate that the retaining heads in APB are robust across different training datasets, demonstrating their generalizability. Consequently, various long-context training data recipes can be employed in the APB framework.
| Training Data | LongAlign | LongAlpaca | AntiHaystack |
| RULER-128K | 81.17 | 81.04 | 80.40 |
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Key R&D Program of China (No.2022ZD0116312) and a grant from the Guoqiang Institute, Tsinghua University. Yuxiang Huang is supported by Beijing National Science Foundation (No. QY24253). We deeply thank the anonymous reviewers for their constructive reviews.
@@ B.1 Training of the Compressor
We use the retaining heads proposed in Huang et al. (2024) as our compressor. Following Huang et al. (2024), we set the intermediate size of the retaining heads to 1024. We utilize the first 3000 samples from the LongAlign (Bai et al., 2024) dataset to generate the training labels, and we train the retaining heads with the a frozen model backbone for 3000 steps in 1 epoch. The batch size set is 1 and the maximum input length is set to 10240. We use AdamW as the optimizer and the learning rate is set to 5e-4, with set to 0.9 and set to 0.95. We apply a linear scheduler whose number of warmup steps is set to 300. Notably, the training loss of the retaining head consists of a regression loss and a smoothing loss, following the setting of Huang et al. (2024), and the balance factor is set to 0.0025. The gradient clipping value is set to 0.5 to avoid gradient explosion.
@@ Appendix D Task Abbreviations
Here, we provide a mapping between the benchmark task abbreviations and their full names used in the main text in Table 10 and 11.
@@ E.2 Hyperparameter Stability
E.3 Short-Context Performance of APB
We report the details and hyperparameters of short-context performance experiment here.
As reperted in Section 4, we compare task performance and inference speed of RULER using FullAttn (implemented with FlashAttn) and APB, with the context length set to 4K tokens, utilizing Llama-3-8B-instruct-1M as the backbone model. In addition, the number of hosts is set to 4, and both anchor length and passing length are set to 256. We evaluate each method on 50 samples per task. We align the settings for inference speed measurement with those used in Figure 3. The detailed results are reported in Table 19.
E.4 Accuracy Experiment on LoCoCo
To present a comparison of APB against optimization strategies that target KV cache size reduction to enable faster inference, we select LoCoCo as a baseline here for comparison with APB. We report its performance in this section. As Table 20 shows, LoCoCo achieves low performance on RULER, as its fixed KV cache size leads to significant information loss, hindering its ability to handle complex long-context tasks effectively. The performance of FullAttn and APB corresponds to the results reported in Table 2 of the main text, which were evaluated using the Llama-3.1-8B-instruct model. Meanwhile, for the evaluation of LoCoCo, we test only 20 samples for each task.
E.5 The Detailed Experimantal Results
Here, we present the detailed experimental results to extensively illustrate each figure in the main text, Section E.1 and E.2. In Section 4.2, the detailed experimental results of Figure 3 are provided in Table 12 and Table 15. Table 12 provides the accurate speed for all methods on Bench, while the speed is the average value across all tasks. By contrast, Table 15 reports the precise speed measurements for all methods on RULER, with the values similarly averaged over all tasks. The performance of each method in Figure 3 is provided in Table 1 and Table 2. In Section 4.3, Figure 17 and Figure 18 present the detailed results corresponding to Figure 4. Specifically, Table 17 aligns with Figure 44 (a) while Table 18 corresponds to Figure 44 (b). In Section 4.4, the detailed experimental results of Figure 5 are provided in Table 16. In Section E.1, we add the breakdown analysis of the prefill and decoding wall-time for various methods, while the results are illustrated in Figure 6. The detailed experimental results are provided in Table 13. In Section 1, the detailed experimental results of Figure 1 are provided in Table 14. The anchor length and passing length are set to 1024 and 512 for 32K input length, 2048 and 1024 for 64K input length, 4096 and 2048 for 128K-1024K input length.